# Video-RAG: Visually-aligned Retrieval-Augmented Long Video Comprehension

### v3

@@ 3.4 Auxiliary Text Retrieval
During the retrieve phase, we employ the Contriever framework to encode the user’s query and the parsed requests for OCR and ASR into text embeddings, then concatenating to form the final query request $\bm{\mathrm{E}}_{req}=\texttt{Contriever}(\texttt{Concat}(\bm{\mathrm{R}},\bm{\mathrm{Q}})),\ \bm{\mathrm{R}}\in\{\bm{\mathrm{R}}_{ocr},\bm{\mathrm{R}}_{asr}\}$. Then we retrieve the auxiliary texts from $DB\in\{DB_{ocr},DB_{asr}\}$ by the FAISS tool, which computes the vector similarity between the query and text chunks stored in the database. Text chunks with a FAISS similarity score greater than threshold $t$ are indexed as the retrieval results $\bm{\mathrm{A}}\in\{\bm{\mathrm{A}}_{ocr},\bm{\mathrm{A}}_{asr}\}$. The process can be formulated as:
[…]
@@
Finally, we acquire the object auxiliary texts based on the object information type retrieval requests $\bm{\mathrm{R}}_{type}$ of the LVLMs in the first phase, which selects and finalizes the object auxiliary information $\bm{\mathrm{A}}_{det}$. $\bm{\mathrm{A}}_{det}$ is one of the elements of the power set $\mathcal{P}$ of $\bm{\mathrm{A}}_{det}^{p}$ selected by $\bm{\mathrm{R}}_{type}$, and the retrieve process can be formulated as:
@@ 4 Experiments
Model #Text LLM Params Frames Short Medium Long Overall Gain
@@
Video-LLaVA [16] - 7B 8 44.6 38.3 35.8 39.6 -
Video-LLaVA + Video-RAG 2.0K 7B 8 49.5 43.0 42.5 45.0 +5.4
LLaVA-NeXT-Video [48] - 7B 16 49.4 43.0 36.7 43.0 -
LLaVA-NeXT-Video + Video-RAG 2.0K 7B 16 56.6 47.4 46.0 50.0 +7.0
LongVA [47] - 7B 32 60.9 49.3 44.0 51.4 -
LongVA + Video-RAG 1.8K 7B 32 65.4 59.1 55.7 60.1 +8.7
Long-LLaVA [43] - 7B 32 60.3 51.4 44.1 52.0 -
Long-LLaVA + Video-RAG 1.9K 7B 32 66.4 60.2 59.8 62.1 +10.1
Qwen2-VL [38] - 72B 32 75.0 63.3 56.3 64.9 -
Qwen2-VL + Video-RAG 2.1K 72B 32 77.4 70.2 71.0 72.9 +8.0
LLaVA-Video [49] - 72B 32 78.0 63.7 59.6 67.1 -
LLaVA-Video + Video-RAG 2.1K 72B 32 81.1 72.9 73.1 75.7 +8.6
Table 1: Performance on the Video-MME [6] benchmark. #Text donates the average token number of auxiliary texts when inferring a single case. By applying our Video-RAG to six LVLMs, we observed an average performance improvement of 8.0% only with the addition of token counts from approximately 14 video frames (144 tokens per frame). In particular, when applying Video-RAG with 72B LLaVA-Video [49], we perform better than the proprietary method Gemini-1.5-Pro [31]. All open-source results are our replication.
4.1 Datasets
Video-MME [6] is a widely used benchmark for assessing the ability of LVLMs to handle detailed videos in real-world scenarios. It is divided into three subsets based on video length, with durations ranging from 11 seconds to 1 hour.
MLVU [52] is a long video understanding benchmark with a large wide of 9 distinct tasks. It is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours with about 12 minutes average video length.
LongVideoBench [41] is a benchmark designed to accurately retrieve and reason over detailed multimodal information from long videos, with 6,678 human-annotated multiple-choice questions in 17 fine-grained categories.
@@ 4.2 Implementation Details
We performed all experiments on NVIDIA A100 80G GPUs. During the auxiliary text generation phase, we first filter the detection requests $\bm{\mathrm{R}}_{det}$ generated by the LVLM to ensure they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. In the auxiliary text retrieval phase, we set both the CLIP similarity threshold and the FAISS similarity threshold $t$ to 0.3. We employ the IndexFlatIP as the similarity calculating method of FAISS [11]. We utilize Long-LLaVA-7B [43] for ablation studies. Since Long-LLaVA is easy to implement while supporting longer context windows, we can investigate the impact of similarity threshold selection in RAG and sampled frame rate on performance. Note that we don’t include the GPT-based Agent methods for comparison due to their resource-intensive nature (complete execution of Video-MME [6] costs around \$2000 for API purchasing when using VideoAgent [4]). Still, we include a mini-experiment of VideoAgent in the Appendix that compares the overall performance, inference time, and GPU requirements with two long-context LVLMs and Video-RAG.
4.3 Main Results
Video-MME. We evaluate our Video-RAG in four 7B open-source LVLMs, including Video-LLaVA [16], LLaVA-NeXT-Video [48], LongVA [47], Long-LLaVA [43], and two 72B LVLM Qwen2-VL [38] and LLaVA-Video [49]. To ensure a fair assessment of the performance improvements introduced by our pipeline, and given the resource constraints (especially 72B LVLMs), we reproduced the performance of all open-source models under the 32-frame setting. Results are shown in Table 1. Specifically, after applying our Video-RAG in 72B LVLM, we perform better than the SOTA proprietary model Gemini-1.5-Pro [31] (75.7% vs. 75.0%). Across the six LVLMs used in our experiments, we gained an average performance boost of 8.0%, especially a significant gain on long videos, demonstrating our pipeline’s effectiveness. This performance improvement is achieved by incorporating token counts from approximately 14 additional video frames (equivalent to 2.0K tokens), each contributing around 144 tokens under most LVLM configurations. We obtain such a large performance enhancement because most LVLMs are pre-trained primarily within the text space and aligned with visual information, often lacking explicit alignment between embedding spaces. Auxiliary texts can serve as semantic supplements sensitive to LVLMs, facilitating model activation and easing the understanding of complex videos.
MLVU. We evaluate Video-RAG when integrating into the 7B and 72B LLaVA-Video [49] of MLVU [52] in a multiple-choice task. As shown in Table 2, when applying Video-RAG in the 7B model, we achieve state-of-the-art performance at scales smaller than 72B using only a 7B model, outperforming the 32B Qryx-1.5 [24] model by 0.1%. Additionally, the 72B LLaVA-Video also has a performance gain of 0.7%. The limited performance gain may be due to having reached a bottleneck of model capacity.
@@ 4.3 Main Results
Video-CCAM [5] 14B 96 63.1
Video-XL [35] 7B 256 64.9
Aria [13] 25.3B 256 70.6
LLaVA-Video* [49] 7B 64 70.8
Oryx-1.5 [24] 32B 128 72.3
LLaVA-Video* [49] 72B 64 73.1
@@
Table 2: The overall performance in the multiple-choice task of the MLVU [52] benchmark. * donates the results of our replication.
LongVideoBench. We evaluate Video-RAG when applied in the 7B and 72B LLaVA-Video [49] of LongVideoBench [41]. We omit the interleaved input format introduced in LongVideoBench when applying Video-RAG. The evaluation results in Table 3 demonstrate that 72B LLaVA-Video with our Video-RAG achieves an overall performance of 65.4% on the validation set. This result surpasses the proprietary LVLM Gemini-1.5-Pro [31] by 1.4%, securing the second place, just 1.3% behind GPT-4o [28]. Meanwhile, the 7B LLaVA-Video also has a performance enhancement of 2.1% when equipped with our Video-RAG.
@@
VideoChat2-Mistral [14] 7B 8 39.3
ShareGPT4Video [2] 7B 8 39.7
LLaVA-Next-Mistral [20] 7B 8 49.1
PLLaVA [42] 34B 16 53.2
LLaVA-Video* [49] 7B 64 56.6
LLaVA-Video* [49] 72B 64 61.9
@@
Table 3: The overall performance on the validation set of LongVideoBench [41]. * donates the results of our replication.
@@ 4.4 Ablation Studies
Effect of different sampling frame number. To explore the effect of the number of sampling frames on Video-RAG, we experience sampling frames number 8, 16, 32, and 64 frames in 7B model Long-LLaVA [43], results are shown in Figure 4. As demonstrated, Video-RAG consistently delivers performance improvements across all frame rates especially in long videos, with these gains increasing as the frame rate rises. Furthermore, the results indicate that the highest accuracy without Video-RAG is achieved when 32 frames are sampled from the video. Therefore, we adopt this configuration in the subsequent ablation experiments.
Figure 4: Performance with different sampling frames rate on Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM.
Effect of different components of Video-RAG. To explore the effectiveness of auxiliary texts, we incrementally add object detection, OCR, and ASR as auxiliary texts after retrieving by the RAG system to evaluate Long-LLaVA-7B [43] in the Video-MME [6] benchmark. As shown in Table 4, the performance of Long-LLaVA progressively improves as auxiliary texts are incrementally added (52.0% → 52.9% → 54.3% → 62.1%). Among these components, ASR auxiliary texts contribute to a general improvement for different video durations, underscoring the critical role of audio transcription that can provide comprehensive information beyond visual cues in video understanding. Meanwhile, we randomly sampled an equivalent token number of auxiliary texts to serve as inputs for assessing the effectiveness of RAG retrieval. The experiment demonstrated a 3.0% improvement in performance after incorporating RAG. We also evaluate the performance across sub-tasks within Video-MME [6] and other video benchmarks like VNBench [51], more details are shown in the Appendix.
Figure 5: Qualitative result shown in Video-MME [6] benchmark when applying Video-RAG with LLaVA-Video [49].
Figure 6: Grad-CAM visualizations of the last hidden state heatmap along with t-SNE visualizations of the user’s query and keyframe features of the example shown in Figure 5. As demonstrated, the retrieved auxiliary texts help cross-modality alignment by assisting the model to pay more attention to query-relevant keyframes and thus generate more robust and accurate answers to the user’s query.
@@
Table 4: Results on combinations of different auxiliary texts in Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM.
Effect of different threshold of RAG processing. When using the RAG tool for retrieval, we specify a similarity threshold $t$ as a criterion for information selection. In the retrieval for OCR and ASR texts, information is selected if its FAISS similarity exceeds $t$. For object detection, frames are selected as keyframes based on their CLIP similarity surpassing $t$, and the relevant information is then extracted. Setting $t$ too high may hinder the retrieval of relevant information while setting it too low can result in information redundancy and increased reasoning complexity. To investigate this trade-off, we conduct ablation experiments to evaluate the impact of different threshold values. The results are shown in Table 5. Notably, $t=0$ and $t=1$ correspond to all auxiliary texts input into the model and no auxiliary texts input, respectively. To balance performance with information density and processing time (especially APE [34] detection in keyframes), we selected a threshold of 0.3 for our implementation. More details about similarity scores are shown in the Appendix. Under this configuration, the additional text length of approximately 1.9K tokens typically remains within the context window limits of open-source LVLMs. For models with more stringent context window limitations, a threshold of 0.4 may also be a viable option.
$t$Token Time Short Medium Long Overall
@@
Table 5: Performance with different thresholds of retrieval on Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM. #Token and Time denote the total token number of the auxiliary texts and the average inference time per question, respectively.

### v4

@@ 3.4 Auxiliary Text Retrieval
During the retrieve phase, we employ Contriever [contriever] to encode the user’s query and the parsed requests for OCR and ASR into text embeddings, then concatenating to form the final query request $\bm{\mathrm{E}}_{req}=\texttt{Contriever}(\texttt{Concat}(\bm{\mathrm{R}},\bm{\mathrm{Q}})),\ \bm{\mathrm{R}}\in\{\bm{\mathrm{R}}_{ocr},\bm{\mathrm{R}}_{asr}\}$. Then we retrieve the auxiliary texts from $DB\in\{DB_{ocr},DB_{asr}\}$ by the FAISS tool, which computes the vector similarity between the query and text chunks stored in the database. Text chunks with a FAISS similarity score greater than threshold $t$ are indexed as the retrieval results $\bm{\mathrm{A}}\in\{\bm{\mathrm{A}}_{ocr},\bm{\mathrm{A}}_{asr}\}$. The process can be formulated as:
$\bm{\mathrm{A}}\xleftarrow{\texttt{Index}}\texttt{FAISS\_similarity}(DB,\bm{\mathrm{E}}_{req})>t$ (9)
Since the text generated by the detection model is in a raw format (“category: [x_min, y_min, length, width]"), it challenges LVLMs to understand the relative relationships between objects. We preprocess the object information using a scene graph, which helps to represent spatial and relational information more explicitly. This preprocessing allows us to construct more coherent and semantically meaningful texts, denoted as $\bm{\mathrm{A}}_{det}^{p}$, which are more readily interpretable by LVLMs. We incorporate three types of object information for each video keyframe: (i) Object Location $\bm{\mathrm{A}}_{loc}$: This refines the positional information of the object, formatted as: “Object {node ID} is a {object category} located at coordinates [x, y] with dimensions {length $\times$ width}” (ii) Object Counting $\bm{\mathrm{A}}_{cnt}$: This counts the number of objects and generates text in the following format: “Object counting: - {object category}: {number}” (iii) Relative Positional Relationships $\bm{\mathrm{A}}_{rel}$: This captures the relative spatial relationships between objects using the format: “Object {node ID} ({object category}) is <positional description> Object {node ID} ({object category})”. By combining this information, we construct a detailed representation of the objects in the frame, denoted as $\bm{\mathrm{A}}_{det}^{p}=\{\bm{\mathrm{A}}_{loc},\bm{\mathrm{A}}_{cnt},\bm{\mathrm{A}}_{rel}\}$:
@@
Finally, we acquire the object auxiliary texts based on the object information type retrieval requests $\bm{\mathrm{R}}_{type}$, which selects and finalizes the object auxiliary information $\bm{\mathrm{A}}_{det}$. $\bm{\mathrm{A}}_{det}$ is one of the elements of the power set $\mathcal{P}$ of $\bm{\mathrm{A}}_{det}^{p}$ selected by $\bm{\mathrm{R}}_{type}$, and the retrieve process can be formulated as:
@@ 4 Experiments
4.1 Datasets
Video-MME [videomme] is a widely used benchmark for assessing the ability of LVLMs to handle detailed videos in real-world scenarios. It is divided into three subsets based on video length, with durations ranging from 11 seconds to 1 hour. MLVU [mlvu] is a long video understanding benchmark with a large wide of 9 distinct tasks. It is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours with about 12 minutes average video length. LongVideoBench [lvb] is a benchmark designed to accurately retrieve and reason over detailed multimodal information from long videos, with 6,678 human-annotated multiple-choice questions in 17 fine-grained categories.
Table 1: Performance on the Video-MME [videomme] benchmark in without subtitles (w/o S), with subtitles (w/ S) and equipped with our Video-RAG (Ours), Frames and Gain means the input frame number and performance gain by applying Video-RAG compared to the baseline with subtitles. By applying Video-RAG to seven LVLMs, we observed an average performance improvement of 2.8% by adding only an average of $\sim$2.0K auxiliary texts compared to $\sim$3.0K full-subtitled tokens per sample. In particular, we perform better when applying Video-RAG with 72B LLaVA-Video [llavavideo] than the proprietary method GPT-4o [gpt4o] (77.4% vs. 77.2%). All results are our republication.
Model Params Frames Short Medium Long Overall Gain
w/o S w/ S Ours w/o S w/ S Ours w/o S w/ S Ours w/o S w/ S Ours
@@ 4.1 Datasets
Video-LLaVA [videollava] 7B 8 45.3 46.1 49.5 38.0 40.7 43.0 36.2 38.1 42.5 39.9 41.6 45.0 +3.4
LLaVA-NeXT-Video [llavanextvideo] 7B 16 49.4 51.8 56.6 43.0 46.4 47.4 36.7 44.9 46.0 43.0 47.7 50.0 +2.3
VITA-1.5 [vita15] 7B 16 67.0 69.9 71.0 54.2 55.7 55.4 47.1 50.4 52.4 56.1 58.7 59.6 +0.9
LongVA [longva] 7B 128 61.1 61.2 66.1 50.4 53.8 60.4 46.2 52.9 59.4 52.6 56.0 62.0 +6.0
Long-LLaVA [longllava] 7B 64 61.9 62.4 67.1 51.4 56.2 60.4 45.4 54.7 60.1 52.9 57.8 62.6 +4.8
Qwen2-VL [qwen2vl] 72B 32 75.0 76.7 77.4 63.3 69.9 70.2 56.3 69.2 71.0 64.9 71.9 72.9 +1.0
LLaVA-Video [llavavideo] 72B 64 80.7 81.8 82.8 68.7 73.8 76.3 62.1 72.2 73.1 70.3 75.9 77.4 +1.5
@@ 4.2 Implementation Details
We performed all experiments on NVIDIA A100 80G GPUs. During the auxiliary text generation phase, we first restrict the detection requests $\bm{\mathrm{R}}_{det}$ generated by LVLMs in decouple prompt then further filter them using spaCy, ensuring they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. In the auxiliary text retrieval phase, we set both the CLIP similarity threshold and the FAISS similarity threshold $t$ to 0.3. We employ the IndexFlatIP as the similarity calculating method of FAISS [faiss]. Note that we don’t include the GPT-based Agent methods for comparison due to their resource-intensive nature (complete execution of Video-MME [videomme] costs around \$2000 for API purchasing when using VideoAgent [videoagent]). Still, we include a mini-experiment of VideoAgent in the Supplemental Material that compares the overall performance, inference time, and GPU requirements with two common long-context LVLMs and our Video-RAG.
Table 2: The overall performance in the multiple-choice task of the MLVU [mlvu] benchmark. * donates the results of our replication.
@@
VITA-1.5 [vita15] 7B 16 60.4
Video-CCAM [videoccam] 14B 96 63.1
Video-XL [videoxl] 7B 256 64.9
Aria [aria] 25.3B 256 70.6
LLaVA-Video* [llavavideo] 7B 64 70.8
Oryx-1.5 [oryx] 32B 128 72.3
LLaVA-Video* [llavavideo] 72B 64 73.1
@@
Table 3: The overall performance on the validation set of LongVideoBench [lvb]. * donates the results of our replication.
@@
VideoChat2-Mistral [videochat] 7B 8 39.3
ShareGPT4Video [chen2024sharegpt4video] 7B 8 39.7
LLaVA-Next-Mistral [llavanext] 7B 8 49.1
PLLaVA [pllava] 34B 16 53.2
VITA-1.5 [vita15] 7B 16 53.6
LLaVA-Video* [llavavideo] 7B 64 56.6
LLaVA-Video* [llavavideo] 72B 64 61.9
@@
4.3 Main Results
Video-MME. We evaluate our Video-RAG in five 7B open-source LVLMs, including Video-LLaVA [videollava], LLaVA-NeXT-Video [llavanextvideo], LongVA [longva], Long-LLaVA [longllava], and two 72B LVLM Qwen2-VL [qwen2vl] and LLaVA-Video [llavavideo]. Constraining by computational resources, we evaluate the LVLMs with their official frame rate setting in Video-MME except for 72B Qwen2-VL, which requires about 3K GPU memory with 768 video frame input ($\sim$38 A100 GPUs). Results are shown in Table 1. Specifically, after applying our Video-RAG in 72B LLaVA-Video [llavavideo], we perform better than the proprietary model GPT-4o [gpt4o] (77.4% vs. 77.2%). Across the seven LVLMs used in our experiments, we gained an average performance boost of 2.8% compared to results with subtitles, especially a significant gain on long videos, demonstrating its effectiveness. This performance improvement is achieved by incorporating token counts from approximately 14 additional video frames (equivalent to 2.0K tokens), each contributing around 144 tokens under most LVLM configurations. We obtain such a large performance enhancement because most LVLMs are pre-trained primarily within the text space and aligned with visual information, often lacking explicit alignment between embedding spaces. Auxiliary texts can serve as semantic supplements sensitive to LVLMs, facilitating model activation and easing the understanding of complex videos.
MLVU. We evaluate Video-RAG when integrating into the 7B and 72B LLaVA-Video [llavavideo] of MLVU [mlvu], a benchmark that is close to performance saturation. As shown in Table 5, Video-RAG’s 1.6% improvement at 7B-scale is substantial, considering that it outperforms the 32B Qryx-1.5 [oryx] by 0.1%, while recent 7B-scale models average only a 1.3% gain (across 15 approaches in MLVU’s leaderboard). Additionally, the 72B LLaVA-Video also has a performance gain of 0.7%, which sets a new state-of-the-art.
LongVideoBench. We evaluate Video-RAG when applied in the 7B and 72B LLaVA-Video [llavavideo] of LongVideoBench [lvb]. We omit the interleaved input format introduced in LongVideoBench when applying Video-RAG. The evaluation results in Table 5 demonstrate that 72B LLaVA-Video with our Video-RAG achieves an overall performance of 65.4% on the validation set. This result surpasses the proprietary LVLM Gemini-1.5-Pro [gemini] by 1.4%, securing the second place, just 1.3% behind GPT-4o [gpt4o]. Meanwhile, the 7B LLaVA-Video also has a performance enhancement of 2.1% when equipped with our Video-RAG.
@@ 4.4 Ablation Studies
Effect of different sampling frame number. To explore the effect of the number of sampling frames on Video-RAG, we experience sampling frames number of 8, 16, 32, 64, 128, and 256 in 7B model LongVA [longva], results are shown in Figure 3. As demonstrated, Video-RAG consistently delivers performance improvements across all frame rates, especially in long videos. The experimental results also indicate that Video-RAG can achieve higher performance gains with fewer frames, demonstrating its potential for applications under resource-constrained conditions.
Figure 3: Performance gain with different sampling frames rate on Video-MME [videomme] when implement LongVA-7B [longllava].
Effect of different components of Video-RAG. To explore the effectiveness of auxiliary texts, we add DET, OCR, and ASR as auxiliary texts before and after retrieving by the RAG to evaluate Long-LLaVA-7B [longllava] with 32-frame setting in the Video-MME [videomme] benchmark. As shown in Table 5, the performance of Long-LLaVA progressively improves as auxiliary texts after retrieving by the RAG system are incrementally added (52.0% → 52.9% → 55.7% → 62.1%). Among these components, ASR auxiliary texts contribute to a general improvement for different video durations, especially for long videos. When all components are integrated, we obtain an optimal performance, as shown in the last row of Table 5. Meanwhile, the experiment shows a 2.3% improvement (59.8% vs 62.1%) in performance after incorporating RAG for retrieval, demonstrating that auxiliary texts after retrieving by the RAG system are query-aligned, which helps cross-modality alignment. We also evaluate across sub-tasks within Video-MME [videomme] and other video benchmarks like MLVU [mlvu], LongVideoBench[lvb], and VNBench [vnbench], more details are shown in the Supplemental Material.
Table 4: Results on combinations of different auxiliary texts in Video-MME [videomme] when using Long-LLaVA-7B [longllava] as the LVLM.
@@
Table 5: Performance with different thresholds of retrieval on Video-MME [videomme] when using Long-LLaVA-7B [longllava] as the LVLM. #Token and Time denote the total token number of the auxiliary texts and the average inference time per question, respectively.
$t$ #Token Time Short Medium Long Overall
@@
rnd 1.9K 11s 65.7 55.8 56.0 59.1
Figure 4: Qualitative result on Video-MME [videomme] when applying Video-RAG with LLaVA-Video [llavavideo].
Figure 5: Grad-CAM visualizations of the last hidden state heatmap along with t-SNE visualizations of the user’s query and keyframe features of the example shown in Figure 4. The retrieved auxiliary texts help cross-modality alignment by assisting the model to pay more attention to query-relevant keyframes and thus generate more robust and accurate answers to the user’s query.
Effect of different thresholds of RAG processing. When retrieving, we specify a similarity threshold $t$ as a criterion for information selection. In the retrieval for OCR and ASR texts, information is selected if its FAISS similarity exceeds $t$. For object detection, frames are selected as keyframes based on their CLIP similarity surpassing $t$, and the relevant information is then extracted. Setting $t$ too high may hinder the retrieval of relevant information while setting it too low can result in information redundancy and increased reasoning complexity. To investigate this trade-off, we conduct ablation experiments to evaluate the impact of different threshold values. The results are shown in Table 5. Notably, $t=0$ and $t=1$ correspond to all auxiliary texts input into the model and no auxiliary texts input, respectively. To balance performance with information density and processing time (especially APE [ape] detection in keyframes), we selected a threshold of 0.3 for our implementation. More details about similarity scores are shown in the Supplemental Material. Under this configuration, the additional text length of approximately 1.9K tokens typically remains within the context window limits of open-source LVLMs. For models with more stringent context window limitations, a threshold of 0.4 may also be a viable option. We also randomly sample an equivalent token number of auxiliary texts to serve as inputs for assessing the effectiveness of RAG retrieval, as shown in the last row of Table 5.